OpenStudy (anonymous):

suppose you are saving money for a new cell phone. at week 0 you have $20 and at week 10 you have $100. if y represents the amount of money you have saved and x represents the number of weeks, find and interpret the slope of the linear equation A. 1/8; $8 is put away each week B. 10; 10$ is put away each week C. 8; $8 is put away each week D 1/10; 10 is put away each week

OpenStudy (anonymous):

I THINK it's C, I asked the same question and still waiting for an answer

OpenStudy (anonymous):

hmm i dont know about that i thought it was that one too but i got 100 at day 9

OpenStudy (anonymous):

think of the linear equation as y=ax + b where x,y are the coordinates, a is the slope, b is the interception of the line to x-axis when x = 0 so, in this question, x is weeks, y is total money, a is slope, and b is what u got at the beginning, $20 y = a (0) + 20= $20 first week now, 100 = a (10) + 20 is the question. you can easily see that a should be 8, that means slope is 8 and you save $8 per week for 10 weeks

OpenStudy (anonymous):

okay i see that but they way i calculated it was i added 8 to each one and i ended up getting 100 as 9 weeks